OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (kamille):

do you know a common linear equation formula?

OpenStudy (gorv):

u have 2 points ..can u calculate the slope ??

OpenStudy (gorv):

for help refer this http://openstudy.com/users/gorv#/updates/543e93b1e4b0c0238da6b6c8

OpenStudy (anonymous):

No, I do not, because my teacher has been on maternity leave and the substitute just kinda sits there not teaching but still giving the work that goes with the lessons.

OpenStudy (anonymous):

I know a really easy way to find the slope that works :)

OpenStudy (anonymous):

without an Equation.. Easier for me anyways

OpenStudy (gorv):

slope =(y2-y1)/(x2-x1)

OpenStudy (gorv):

take one point (x1,y1) and other as (x2,y2)

OpenStudy (anonymous):

@*Shadow* Oh?

OpenStudy (gorv):

will u try ??

OpenStudy (anonymous):

Yes, but as I said I don't even know how or what to do and I am not familiar with any of this vocabulary due to poor education.

OpenStudy (anonymous):

Ok, fukaawa.

OpenStudy (anonymous):

My way is. find the difference between both (x) coordinates.

OpenStudy (anonymous):

@fukaawa

OpenStudy (academicgurusinc):

Math can be simple if the steps are broken down and you try your hardest. We understand that you are facing some limitation so try and follow along with the steps below. Answer: Linear formula: y=mx + b m= slope =(change in y)/(change in x) b= y- intercept Step 1: calculate slope m= slope =(change in y)/(change in x)= [2 - (-4)]/ [-2-0]=6/-2=-3 Step 2: calculate y-intercept by plugging in point values and slope value. ex (-2, 2) y= mx + b 2 = -3 (-2) + b 2 =6 + b 2-6 = b -4 =b Step 3: rewrite equation with slope and y-intercept values y= -3 x -4 Was this helpful? If you need additional clarification, feel free to drop us a line at info@academicgurusinc.com We hope we were able to help. Follow us on Facebook (AcademicGurusInc) and twitter (@Academic_Gurus).